solve for x: x + 5 = 4 square root of x + 1
Answer by edjones(8007)

x+5=4sqrt(x+1)
x^2+10x+25=16(x+1) square each side.
x^2+10x+25=16x+16
x^2-6x+9=0
(x-3)^2=0
x=3
.
Ed